OpenStudy (anonymous):

integrate 2cos(pix/2) this is entirely weird.

OpenStudy (anonymous):

@undeadknight26 any idea? :3

undeadknight26 (undeadknight26):

I no even learneded this yet o.o

OpenStudy (anonymous):

crap okay d'you know anyone here who is familiar with integration? ^^

OpenStudy (anonymous):

@zepdrix ?

zepdrix (zepdrix):

Let's write it like this: \[\Large\rm \int 2\cos\left(\frac{\pi}{2}x\right)dx\]What's weird about it? XD Shouldn't be too bad.

OpenStudy (anonymous):

well, i did that and i got 2(2/pi)sin(pi/2)x +c

zepdrix (zepdrix):

Mmmm ya looks great! \c:/

OpenStudy (anonymous):

oh cool okay thanks